1.) The physcian ordered 1/2 NS to run over 6 hours with a drip factor of 20gtt/mL. Calculate the correct IV flow rate to set on the infusion pump. Round to the nearest whole number.

2.) The physican ordered 750 mL D5 NS to infuse over 6 hours with a drop factor of 10 gtt/mL. Caculate the correct IV flow rate in mL per hour.

Round to nearest whole number.
Answer #1

Q.1

1/2 NS= 500ml

Over = 6 hours

Drop factor= 20

Formula= Amount of fluid *drop factor/ time*min

= 500*20/6×60

= 27.77 drops/min

That is 28drops/min.

Q.2:

Order

D5 NS= 750ml

Time 6 hours

Drop factor 10

Formula= amount of fluid×drop factor/time×min

= 750×10/ 6×60

= 20.8 drops/ min

That is

21 drops /min.